OpenStudy (anonymous):

find roots x^4-4x^3+7x^2-4x+6=0

OpenStudy (jamesj):

Try +1, -1, +2, -2; and if none of them work, try +i, -i.

OpenStudy (jamesj):

Notice btw that if +i is a root, then -i must also be a root because all the coefficients of the polynomial are real. That being the case, you can now factor out a quadratic and the problem becomes easy.

OpenStudy (anonymous):

can we use synthetic division with i???

OpenStudy (anonymous):

reply please.........

OpenStudy (anonymous):

i is the root. i have applied synthetic division and get a very complicated ans i.e (x-i)(x^3+(-4+i)x^2+(6-4i)x+6i)=0 how can i do further?

OpenStudy (jamesj):

If i is a root, then -i is a root and hence \[ (x+i)(x=-i) = x^2 + 1 \] is a divisor of the polynomial; i.e., you can divide the polynomial by \( x^2 + 1 \).

OpenStudy (jamesj):

\[ (x+1)(x-i) = x^2 + 1 \]

OpenStudy (anonymous):

my method is correct?

OpenStudy (jamesj):

Probably, but unduly complicated. Remember, if a polynomial has all real coefficients, and a complex number \( a \) is a root, then the complex conjugate of \( a \) is also a root and hence \( x^2 + |a|^2 \) is always a factor.

OpenStudy (jamesj):

If you factor out \( x^2 + 1\) you will find a much simpler polynomial to deal with: just a quadratic.

OpenStudy (anonymous):

(x+1)(x−i)=x^2+1 how do you do this step?

OpenStudy (anonymous):

please explain if u can?

OpenStudy (jamesj):

In general, \( (a-b)(a+b) = a^2 - b^2 \) thus \[ (x+i)(x-i) = x^2 - i^2 = x^2 -(-1) = x^2 + 1 \]